OpenStudy (anonymous):

If 7.400 g of C6H6 is burned and the heat produced from the burning is added to 5691 g of water at 21 °C, what is the final temperature of the water?

OpenStudy (anonymous):

the kj to the c6h6 is 6542 and there are 2 moles it as well. For some reason i can't seem to get the right number.

OpenStudy (aaronq):

you have to use: \(q=m*C_p*\Delta T\) don't forget that \(\Delta T =T_f-T_i\) so the equation is: \(q=m*C_p*(T_f-T_i)\)

OpenStudy (anonymous):

you're a really helpful, but im just an idiot and used the wrong molar mass by accident.

OpenStudy (aaronq):

haha ah damn